explain bragg equation.
1
Expert's answer
2020-09-28T08:11:32-0400

Solution

According to bragg, if any light with

"\lambda" wavelength incident on nth plane of any crystal (which has d is inter planer distance) with angle "\theta" then bragg equation can be written as below

"2d\sin\theta=n\lambda"